# AdaSVD: Adaptive Singular Value Decomposition for Large Language Models

### v1

@@ Abstract
Large language models (LLMs) have achieved remarkable success in natural language processing (NLP) tasks, yet their substantial memory requirements present significant challenges for deployment on resource-constrained devices. Singular Value Decomposition (SVD) has emerged as a promising compression technique for LLMs, offering considerable reductions in memory overhead. However, existing SVD-based methods often struggle to effectively mitigate the errors introduced by SVD truncation, leading to a noticeable performance gap when compared to the original models. Furthermore, applying a uniform compression ratio across all transformer layers fails to account for the varying importance of different layers. To address these challenges, we propose AdaSVD, an adaptive SVD-based LLM compression approach. Specifically, AdaSVD introduces adaComp, which adaptively compensates for SVD truncation errors by alternately updating the singular matrices and . Additionally, AdaSVD introduces adaCR, which adaptively assigns layer-specific compression ratios based on the relative importance of each layer. Extensive experiments across multiple LLM families and evaluation metrics demonstrate that AdaSVD consistently outperforms state-of-the-art (SOTA) SVD-based methods, achieving superior performance with significantly reduced memory requirements. The code and models will be available at https://github.com/ZHITENGLI/AdaSVD.
@@ 1 Introduction
Recently, large language models (LLMs) based on the Transformer architecture (Vaswani, 2017) have shown remarkable potential across a wide range of natural language processing (NLP) tasks. However, their success is largely driven by their massive scale, with models such as the LLaMA family (Touvron et al., 2023a) and the Open Pre-trained Transformer (OPT) series (Zhang et al., 2022) containing up to 70B and 66B parameters, respectively. The substantial memory requirements of these models present significant challenges for deploying them on mobile devices. Consequently, the widespread adoption of LLMs remains limited by their immense resource demands (Wan et al., 2023; Wang et al., 2024a; Zhou et al., 2024).
Recent research on large language model (LLM) compression has explored various techniques, including weight quantization (Lin et al., 2024; Frantar et al., 2023), network pruning (Sun et al., 2024; Frantar & Alistarh, 2023), low-rank factorization (Wang et al., 2024b; Zhang et al., 2024; Yuan et al., 2024), and knowledge distillation (Zhong et al., 2024; Gu et al., 2024). Among these methods, low-rank factorization using Singular Value Decomposition (SVD) (Hsu et al., 2022a; Yuan et al., 2024; Wang et al., 2024b) stands out as a powerful approach for reducing both model size and computational cost. SVD achieves this by decomposing large weight matrices into smaller, low-rank components while preserving model performance. Since LLMs are often memory-bound during inference (Dao et al., 2022; Dao, 2024), SVD compression can effectively accelerate model inference by reducing the memory requirements, even when applied solely to the weights. This approach does not require specialized hardware or custom operators, unlike weight quantization, making SVD more versatile across different platforms. Additionally, SVD is orthogonal to other compression techniques (Wang et al., 2024b), allowing it to be combined with methods like weight quantization or network pruning for even greater efficiency, enabling more scalable and adaptable solutions for deploying LLMs.
Recent advancements in SVD-based LLM compression, including FWSVD (Hsu et al., 2022a), ASVD (Yuan et al., 2024), and SVD-LLM (Wang et al., 2024b), have significantly improved the low-rank factorization approach, enhancing the overall effectiveness of SVD compression. For example, FWSVD introduces Fisher information to prioritize the importance of parameters, while ASVD accounts for the impact of activation distribution on compression error. SVD-LLM establishes a relationship between singular values and compression loss through the data whitening techniques. While these methods have led to notable improvements in SVD compression, they still face challenges when applied at high compression ratios.
@@
We propose adaCR, an adaptive compression ratio method that assigns layer-specific compression ratios based on their relative importance for LLMs. This importance-aware approach outperforms the previously adopted uniform compression ratio.
Extensive experiments demonstrate that our method, AdaSVD, significantly outperforms the previous SOTA SVD-based LLM compression method, SVD-LLM, effectively narrowing the performance gap between compressed and original models.
@@ 2.1 LLM Compression Techniques
Recent advancements in model compression techniques have significantly enhanced the efficiency of deploying LLMs while maintaining their performance. Widely explored approaches include weight quantization (Frantar et al., 2023; Lin et al., 2024), network pruning (Frantar & Alistarh, 2023; Ma et al., 2023; Yang et al., 2024; Gromov et al., 2024; Ashkboos et al., 2024), and hybrid methods (Dong et al., 2025). In unstructured pruning, SparseGPT (Frantar & Alistarh, 2023) prunes weights based on their importance, as determined by the Hessian matrix. However, it faces challenges in achieving optimal speedup, particularly due to hardware compatibility issues. Structured pruning methods, in contrast, are more hardware-friendly. LLM-Pruner (Ma et al., 2023) selectively removes non-critical coupled structures using gradient information. LaCo (Yang et al., 2024) introduces a layer-wise pruning strategy, where subsequent layers collapse into preceding ones. Gromov et al. (2024) explores the effectiveness of basic layer-pruning techniques combined with parameter-efficient fine-tuning (PEFT). Additionally, SliceGPT (Ashkboos et al., 2024) has pioneered post-training sparsification, emphasizing the importance of layer removal order for optimal performance. Quantization techniques offer another significant avenue for compression. GPTQ (Frantar et al., 2023) applies layer-wise quantization and reduces quantization errors through second-order error compensation. AWQ (Lin et al., 2024) introduces activation-aware weight quantization, employing a scale transformation between weights and activations. Moreover, BiLLM (Huang et al., 2024) and ARB-LLM (Li et al., 2025) achieve further compression to 1-bit while maintaining remarkable performance. More recently, STB-LLM (Dong et al., 2025) combines 1-bit quantization with pruning to achieve even greater memory reduction for LLMs. However, many of these compression techniques face challenges related to hardware compatibility, often requiring custom CUDA kernels (Dong et al., 2025) to enable real-time inference speedup.
@@ 2.2 SVD-based LLM Compression
Singular Value Decomposition (SVD) is a widely used technique for reducing matrix size by approximating a matrix with two smaller, low-rank matrices (Golub et al., 1987). Although SVD-based methods have demonstrated potential in compressing LLMs, their full capabilities remain underexplored. Standard SVD typically focuses on compressing the original weight matrix without considering the significance of individual parameters, which can lead to considerable compression errors. To address this, Hsu et al. (2022b) introduced FWSVD, which incorporates Fisher information to weight the importance of parameters. However, this method requires complex gradient calculations, making it resource-intensive. Another limitation of standard SVD is the impact of activation distribution on compression errors. To mitigate this, Yuan et al. (2024) proposed ASVD, which scales the weight matrix with a diagonal matrix that accounts for the influence of input channels on the weights. Subsequently, Wang et al. (2024b) introduced SVD-LLM, which establishes a connection between singular values and compression loss. This work demonstrates that truncating the smallest singular values after data whitening effectively minimizes compression loss. Despite these advancements, existing methods still exhibit significant accuracy loss at higher compression ratios and lack a comprehensive approach for compensating compressed weights after SVD truncation. Furthermore, most methods apply a uniform compression ratio across all transformer layers, overlooking the varying importance of different layers. Our AdaSVD seeks to address these limitations by proposing an adaptive compensation method (adaComp) and an importance-aware adaptive compression ratio method (adaCR).
@@ 3 Method
Overview. As illustrated in Figure 2, our AdaSVD integrates adaptive compensation for SVD truncation (adaComp) with an adaptive importance-aware compression ratio method (adaCR). In Section 3.1, we first describe how adaComp compensates for SVD truncation. Next, in Section 3.2, we detail how adaCR determines the compression ratio based on layer importance. The pseudocode of AdaSVD is shown in Algorithm 1, and pseudocodes for adaComp and adaCR are provided in supplementary file.
@@ 3.1 Adaptive Compensation for SVD Truncation
Previous works (Hsu et al., 2022b; Yuan et al., 2024; Wang et al., 2024b) have made significant efforts to minimize . However, some of these methods involve complex and time-consuming preprocessing steps. Furthermore, they still face substantial challenges in effectively mitigating the large errors that arise under high compression ratios, particularly when truncating 50% or more of the parameters.
@@
and then obtain the solution for by using the Moore-Penrose pseudoinverse (Penrose, 1955) of :
@@
where , and cardinality .
As shown in Figure 3 (b), integrating stack-of-batch strategy further reduces the compression error.
@@
where and denote the updated singular matrices after -th iteration, respectively. As shown in Figure 3 (c), the gap between the outputs of the compressed and original models narrows after alternative updates. The overlapping area rapidly increases after just a few iterations. More visual comparisons are shown in supplementary file.
@@ 3.2 Adaptive SVD Compression Ratio
Previous studies on SVD compression typically apply a uniform compression ratio across all transformer layers of LLMs, overlooking the varying importance of different layers. Inspired by Men et al. (2024) and Dumitru et al. (2024), we propose adaCR, which adaptively determines the SVD compression ratio for each layer, considering each layer’s impact on activations.
@@
where denotes the importance of . The similarity metric used can vary, and for simplicity, we adopt cosine similarity in our method.
Then, we normalize by mean centering to obtain the relative importance:
@@
Given the compression ratio for the -th layer, we truncate the vectors of largest singular values from both and so that
@@
| Ratio | Method | WikiText-2 | PTB | C4 | Openb. | ARC_e | WinoG. | HellaS. | ARC_c | PIQA | Boolq | Average |
| 0% | Original | 5.68 | 8.35 | 7.34 | 0.28 | 0.67 | 0.67 | 0.56 | 0.38 | 0.78 | 0.27 | 0.52 |
| 40% | SVD | 39,661.00 | 69,493.00 | 56,954.00 | 0.14 | 0.26 | 0.49 | 0.26 | 0.21 | 0.53 | 0.46 | 0.34 |
| SVD-LLM | 16.11 | 719.44 | 61.95 | 0.17 | 0.37 | 0.56 | 0.30 | 0.21 | 0.57 | 0.39 | 0.37 | |
| AdaSVD | 14.76 | 304.62 | 56.98 | 0.19 | 0.41 | 0.58 | 0.32 | 0.23 | 0.58 | 0.39 | 0.39 | |
| 50% | SVD | 53,999.00 | 39,207.00 | 58,558.00 | 0.14 | 0.26 | 0.47 | 0.26 | 0.22 | 0.53 | 0.39 | 0.32 |
| SVD-LLM | 27.19 | 1,772.91 | 129.66 | 0.14 | 0.32 | 0.51 | 0.28 | 0.18 | 0.54 | 0.38 | 0.34 | |
| AdaSVD | 25.58 | 593.14 | 113.84 | 0.15 | 0.34 | 0.54 | 0.29 | 0.20 | 0.56 | 0.38 | 0.35 | |
| 60% | SVD | 65,186.00 | 79,164.00 | 70,381.00 | 0.15 | 0.24 | 0.52 | 0.25 | 0.23 | 0.53 | 0.50 | 0.35 |
| SVD-LLM | 89.90 | 2,052.89 | 561.00 | 0.13 | 0.26 | 0.47 | 0.27 | 0.19 | 0.53 | 0.37 | 0.32 | |
| AdaSVD | 60.08 | 2,137.28 | 294.26 | 0.12 | 0.27 | 0.50 | 0.27 | 0.20 | 0.53 | 0.38 | 0.32 | |
| 70% | SVD | 40,011.00 | 333,773.00 | 55,284.00 | 0.16 | 0.26 | 0.50 | 0.25 | 0.23 | 0.53 | 0.39 | 0.33 |
| SVD-LLM | 125.16 | 6,139.78 | 677.38 | 0.14 | 0.26 | 0.49 | 0.26 | 0.21 | 0.53 | 0.38 | 0.32 | |
| AdaSVD | 107.90 | 5,027.62 | 441.33 | 0.12 | 0.26 | 0.48 | 0.27 | 0.20 | 0.53 | 0.38 | 0.32 | |
| 80% | SVD | 24,965.00 | 41,571.00 | 53,894.00 | 0.15 | 0.26 | 0.48 | 0.26 | 0.23 | 0.51 | 0.41 | 0.33 |
| SVD-LLM | 372.48 | 6,268.53 | 1,688.78 | 0.14 | 0.26 | 0.49 | 0.26 | 0.23 | 0.52 | 0.38 | 0.33 | |
| AdaSVD | 206.51 | 6,613.44 | 679.66 | 0.12 | 0.26 | 0.49 | 0.26 | 0.20 | 0.53 | 0.38 | 0.32 |
@@ 4.1 Setup
We compare the proposed AdaSVD against four baselines, including vanilla SVD and SOTA SVD-based LLM compression methods FWSVD (Hsu et al., 2022b), ASVD (Yuan et al., 2024) and SVD-LLM (Wang et al., 2024b).
Models and Datasets. To demonstrate the generalizability of our method, we evaluate the performance of AdaSVD and the baselines on five models from three different LLM families, including LLaMA-7B (Touvron et al., 2023a), LLaMA2-7B (Touvron et al., 2023b), OPT-6.7B (Zhang et al., 2022), Mistral (Jiang et al., 2023), and Vicuna-7B (Chiang et al., 2023). We also use 10 datasets, including three language modeling datasets (WikiText-2 (Merity et al., 2016), PTB (Marcus et al., 1993), and C4 (Raffel et al., 2023)) and seven common-sense reasoning datasets (OpenbookQA (Mihaylov et al., 2018), WinoGrande (Sakaguchi et al., 2019), HellaSwag (Zellers et al., 2019), PIQA (Bisk et al., 2019), BoolQ (Clark et al., 2019), ARC-e, and ARC-c (Clark et al., 2018)). We use the LM-Evaluation-Harness framework (Gao et al., 2023) to evaluate the model performance on these zero-shot QA datasets.
Implementation Details. To ensure a fair comparison, we followed ASVD (Yuan et al., 2024) and SVD-LLM (Wang et al., 2024b) to randomly select 256 samples from WikiText-2 as the calibration data and conduct data whitening before SVD truncation. All the experiments are conducted with PyTorch (Paszke et al., 2019b) and Huggingface (Paszke et al., 2019a) on a single NVIDIA A100-80GB GPU.

### v3

@@ Abstract
Large language models (LLMs) have achieved remarkable success in natural language processing (NLP) tasks, yet their substantial memory requirements present significant challenges for deployment on resource-constrained devices. Singular Value Decomposition (SVD) has emerged as a promising compression technique for LLMs, offering considerable reductions in memory overhead. However, existing SVD-based methods often struggle to effectively mitigate the errors introduced by SVD truncation, leading to a noticeable performance gap when compared to the original models. Furthermore, applying a uniform compression ratio across all transformer layers fails to account for the varying importance of different layers. To address these challenges, we propose AdaSVD, an adaptive SVD-based LLM compression approach. Specifically, AdaSVD introduces adaComp, which adaptively compensates for SVD truncation errors by alternately updating the singular matrices and . Additionally, AdaSVD introduces adaCR, which adaptively assigns layer-specific compression ratios based on the relative importance of each layer. Extensive experiments across multiple LLM/VLM families and evaluation metrics demonstrate that AdaSVD consistently outperforms state-of-the-art (SOTA) SVD-based methods, achieving superior performance with significantly reduced memory requirements. Code and models of AdaSVD will be available at https://github.com/ZHITENGLI/AdaSVD.
@@ 1 Introduction
Recently, large language models (LLMs) based on the Transformer architecture [37] have shown remarkable potential across a wide range of natural language processing (NLP) tasks. However, their success is largely driven by their massive scale, with models such as the LLaMA family [35] and the Open Pre-trained Transformer (OPT) series [45] containing up to 70B and 66B parameters, respectively. The substantial memory requirements of these models present significant challenges for deploying them on mobile devices. Consequently, the widespread adoption of LLMs remains limited by their immense resource demands [38, 39, 47].
Recent research on large language model (LLM) compression has explored various techniques, including weight quantization [22, 11], network pruning [34, 10], low-rank factorization [40, 44, 42], and knowledge distillation [46, 15]. Among these methods, low-rank factorization using Singular Value Decomposition (SVD) [17, 42, 40] stands out as a powerful approach for reducing both model size and computational cost. SVD achieves this by decomposing large weight matrices into smaller, low-rank components while preserving model performance. Since LLMs are often memory-bound during inference [7, 6], SVD compression can effectively accelerate model inference by reducing the memory requirements, even when applied solely to the weights. This approach does not require specialized hardware or custom operators, unlike weight quantization, making SVD more versatile across different platforms. Additionally, SVD is orthogonal to other compression techniques [40], allowing it to be combined with methods like weight quantization or network pruning for even greater efficiency, enabling more scalable and adaptable solutions for deploying LLMs.
Recent advancements in SVD-based LLM compression, including FWSVD [17], ASVD [42], and SVD-LLM [40], have significantly improved the low-rank factorization approach, enhancing the overall effectiveness of SVD compression. For example, FWSVD introduces Fisher information to prioritize the importance of parameters, while ASVD accounts for the impact of activation distribution on compression error. SVD-LLM establishes a relationship between singular values and compression loss through the data whitening techniques. While these methods have led to notable improvements in SVD compression, they still face challenges when applied at high compression ratios.
@@
We propose adaCR, an adaptive compression ratio method that assigns layer-specific compression ratios according to their relative importance in LLMs. This importance-aware approach outperforms the previously used uniform compression ratio method.
Extensive experiments on LLMs/VLMs demonstrate that our method, AdaSVD, significantly outperforms the previous SOTA SVD-based LLM compression method, SVD-LLM, effectively narrowing the performance gap between compressed and original models.
@@ 2.1 LLM Compression Techniques
Recent advancements in model compression techniques have significantly enhanced the efficiency of deploying LLMs while maintaining their performance. Widely explored approaches include weight quantization [11, 22], network pruning [10, 25, 41, 14, 1], and hybrid methods [8]. In unstructured pruning, SparseGPT [10] prunes weights based on their importance, as determined by the Hessian matrix. However, it faces challenges in achieving optimal speedup, particularly due to hardware compatibility issues. Structured pruning methods, in contrast, are more hardware-friendly. LLM-Pruner [25] selectively removes non-critical coupled structures using gradient information. LaCo [41] introduces a layer-wise pruning strategy, where subsequent layers collapse into preceding ones. Gromov et al. [14] explores the effectiveness of basic layer-pruning techniques combined with parameter-efficient fine-tuning (PEFT). Additionally, SliceGPT [1] has pioneered post-training sparsification, emphasizing the importance of layer removal order for optimal performance. Quantization techniques offer another significant avenue for compression. GPTQ [11] applies layer-wise quantization and reduces quantization errors through second-order error compensation. AWQ [22] introduces activation-aware weight quantization, employing a scale transformation between weights and activations. Moreover, BiLLM [19] and ARB-LLM [21] achieve further compression to 1-bit while maintaining remarkable performance. More recently, STB-LLM [8] combines 1-bit quantization with pruning to achieve even greater memory reduction for LLMs. However, many of these compression techniques face challenges related to hardware compatibility, often requiring custom CUDA kernels [8] to enable real-time inference speedup.
@@ 2.2 SVD-based LLM Compression
Singular Value Decomposition (SVD) is a widely used technique for reducing matrix size by approximating a matrix with two smaller, low-rank matrices [13]. Although SVD-based methods have demonstrated potential in compressing LLMs, their full capabilities remain underexplored. Standard SVD typically focuses on compressing the original weight matrix without considering the significance of individual parameters, which can lead to considerable compression errors. To address this, Hsu et al. [18] introduced FWSVD, which incorporates Fisher information to weight the importance of parameters. However, this method requires complex gradient calculations, making it resource-intensive. Another limitation of standard SVD is the impact of activation distribution on compression errors. To mitigate this, Yuan et al. [42] proposed ASVD, which scales the weight matrix with a diagonal matrix that accounts for the influence of input channels on the weights. Subsequently, Wang et al. [40] introduced SVD-LLM, which establishes a connection between singular values and compression loss. This work demonstrates that truncating the smallest singular values after data whitening effectively minimizes compression loss. Despite these advancements, existing methods still exhibit significant accuracy loss at higher compression ratios and lack a comprehensive approach for compensating compressed weights after SVD truncation. Furthermore, most methods apply a uniform compression ratio across all transformer layers, overlooking the varying importance of different layers. AdaSVD seeks to address these limitations by proposing an adaptive compensation method (adaComp) and an importance-aware adaptive compression ratio method (adaCR).
@@ 3 Method
Overview. As illustrated in Figure 2, our AdaSVD integrates adaptive compensation for SVD truncation (adaComp) with an adaptive importance-aware compression ratio method (adaCR). In Sec. 3.1, we first describe how adaComp compensates for SVD truncation. Next, in Sec. 3.2, we detail how adaCR determines the compression ratio based on layer importance. The pseudocode of AdaSVD is shown in Algorithm 1, and pseudocodes for adaComp and adaCR are provided in the supplementary file.
@@ 3.1 Adaptive Compensation for SVD Truncation
Previous works [18, 42, 40] have made significant efforts to minimize . However, some of them involve complex and time-consuming preprocessing steps. Furthermore, they still face substantial challenges in effectively mitigating the large errors that arise under high compression ratios, particularly when truncating 60% or more of the parameters.
@@
and then obtain the solution for by using the Moore-Penrose pseudoinverse [31] of :
@@
where , and cardinality . As shown in Figure 3 (b), integrating the stack-of-batch strategy further reduces the compression error.
@@
where and denote the updated singular matrices after -th iteration, respectively, while the region bounded by corresponding to one iteration of alternative update. As shown in Figure 3 (c), the gap between the outputs of the compressed and original models narrows after alternative updates. The overlapping area rapidly increases after just a few iterations. More visual comparisons are shown in the supplementary file.
Notably, our adaptive compensation can be integrated with data whitening proposed by Wang et al. [40] and Liu et al. [24], further reducing the SVD truncation error.
Ratio | Method | WikiText-2 | PTB | C4 | Mmlu | ARC_e | WinoG. | HellaS. | PIQA | Average |
0% | Original | 5.68 | 8.35 | 7.34 | 45.30 | 74.62 | 69.22 | 76.00 | 79.11 | 68.85 |
40% | SVD | 39,661.03 | 69,493.00 | 56,954.00 | 26.51 | 26.39 | 48.62 | 25.64 | 52.99 | 36.03 |
FWSVD [18] | 8,060.35 | 9,684.10 | 7,955.21 | 25.74 | 26.05 | 50.20 | 25.70 | 52.39 | 36.01 | |
ASVD [42] | 1,609.32 | 7,319.49 | 1,271.85 | 24.35 | 26.81 | 49.49 | 25.83 | 53.81 | 36.06 | |
SVD-LLM [40] | 16.11 | 719.44 | 61.95 | 22.97 | 36.99 | 56.04 | 30.49 | 56.96 | 40.69 | |
AdaSVD | 14.76 (8%) | 304.62 (58%) | 56.98 (8%) | 23.63 | 41.12 | 58.17 | 31.75 | 58.49 | 42.63 | |
50% | SVD | 53,999.48 | 39,207.00 | 58,558.00 | 25.43 | 25.80 | 47.36 | 25.55 | 52.67 | 35.36 |
FWSVD [18] | 8,173.21 | 8,615.71 | 8,024.67 | 24.83 | 25.84 | 48.70 | 25.64 | 52.83 | 35.57 | |
ASVD [42] | 6,977.57 | 15,539.44 | 4,785.15 | 24.52 | 25.13 | 49.17 | 25.48 | 52.94 | 35.45 | |
SVD-LLM [40] | 27.19 | 1,772.91 | 129.66 | 23.44 | 31.65 | 51.14 | 28.38 | 54.57 | 37.83 | |
AdaSVD | 25.58 (6%) | 593.14 (67%) | 113.84 (12%) | 23.24 | 34.18 | 54.06 | 28.88 | 55.50 | 39.17 | |
60% | SVD | 65,186.67 | 79,164.00 | 70,381.00 | 22.94 | 24.49 | 51.85 | 25.40 | 53.16 | 35.57 |
FWSVD [18] | 27,213.30 | 24,962.80 | 47,284.87 | 26.91 | 25.38 | 48.46 | 25.61 | 51.96 | 35.66 | |
ASVD [42] | 10,003.57 | 15,530.19 | 9,983.83 | 26.89 | 26.68 | 48.86 | 25.76 | 51.80 | 36.00 | |
SVD-LLM [40] | 89.90 | 2,052.89 | 561.00 | 22.88 | 26.73 | 47.43 | 26.89 | 53.48 | 35.48 | |
AdaSVD | 50.33 (44%) | 1,216.95 (41%) | 239.18 (57%) | 24.69 | 28.20 | 51.22 | 27.36 | 52.83 | 36.87 |
@@ 3.2 Adaptive SVD Compression Ratio
Previous studies on SVD compression typically apply a uniform compression ratio across all transformer layers of LLMs, overlooking the varying importance of different layers. Inspired by Men et al. [27] and Dumitru et al. [9], we propose adaCR, which adaptively determines the SVD compression ratio for each transformer layer, considering each layer’s distinct impact on activations.
@@
where denotes the layer-wise importance of . The similarity metric used can vary, and for simplicity, we adopt cosine similarity in our method.
Then, we normalize through mean centering to obtain the relative importance of :
@@
Given the compression ratio for the -th layer by adaCR, we truncate the vectors of least singular values from both and so that
@@ 4.1 Setup
We compare our AdaSVD with four baselines, including vanilla SVD and SOTA SVD-based LLM compression methods FWSVD [18], ASVD [42], and SVD-LLM [40].
Method | OPT-6.7B | LLaMA2-7B | Mistral-7B | Vicuna-7B |
SVD | 18,607.24 | 65,186.67 | 30,378.35 | 78,704.50 |
FWSVD [18] | 8,569.56 | 27,213.30 | 5,481.24 | 8,185.66 |
ASVD [42] | 10,326.48 | 10,003.57 | 22,705.51 | 20,241.17 |
SVD-LLM [40] | 92.10 | 89.90 | 72.17 | 64.06 |
AdaSVD | 86.64 (6%) | 50.33 (44%) | 67.22 (7%) | 56.97 (11%) |
Models and Datasets. To demonstrate the generalizability of our method, we evaluate the performance of AdaSVD and the baselines on four models from three different LLM families, including LLaMA2-7B [36], OPT-6.7B [45], Mistral-7B [20], and Vicuna-7B [4]. We benchmark on eight datasets, including three language modeling datasets (WikiText-2 [28], PTB [26], and C4 [32]) and five common-sense reasoning datasets (WinoGrande [33], HellaSwag [43], PIQA [2], ARC-e [5], and Mmlu [16]). We use the LM-Evaluation-Harness framework [12] to evaluate the model performance on these zero-shot Question-Answering (QA) datasets.
Implementation Details. To ensure a fair comparison, we followed ASVD [42] and SVD-LLM [40] to randomly select 256 samples from WikiText-2 as the calibration data and conduct data whitening before SVD truncation. All the experiments are conducted with PyTorch [30] and Huggingface [29] on a single NVIDIA A100-80GB GPU.